# Constraining the Nuclear Equation of State of neutron star via high-frequency quasi-periodic oscillation in short gamma-ray bursts

### v1

@@ Abstract
The determination of the equation of state (EOS) of a neutron star (NS) and its maximum mass is very important for understanding the formation and properties of NSs under extreme conditions, but they remain open questions. Short-duration gamma-ray bursts (GRBs) are believed to originate from the merger of binary NSs or giant flares (GFs) of soft gamma-ray repeaters (SGRs). Recently, the high-frequency quasi-periodic oscillations (QPOs) have been claimed to be identified from two short GRBs (931101B and 910711). In this paper, we propose that the observed high-frequency QPOs in these two short GRBs result from torsional oscillations in the GFs of SGRs associated with cold NSs, or from radial oscillations of hypermassive NSs (HMNSs) as the hot remnants of binary NS mergers, and then to constrain the EOS of NSs. For torsional oscillations, the six selected EOSs (TM1, NL3, APR, SLy4, DDME2, and GM1) of NSs suitable for the zero-temperature condition exhibit significant overlap in mass ranges, suggesting that we cannot constrain the EOS of NSs. For radial oscillations, the six selected EOSs (IUF, TM1, TMA, FSG, BHBLp, and NL3) of NSs suitable for the high-temperature condition cannot be ruled out when redshift is considered. However, it is found that the EOS can only be constrained if the redshift and temperature of the remnant can be measured.
@@ 1 Introduction
Neutron star (NS) is one of the most intriguing compact objects in astrophysical studies due to its extreme properties. The maximum mass and equation of state (EOS) of NSs play an important role in understanding their structure and properties (Shapiro & Teukolsky, 1983). The first estimate of the maximum mass of a NS ($M_{\text{max}}$) for a static equilibrium solution in early study was $M_{\text{max}}\sim\,0.72\ M_{\odot}$ by simplifying the NS as a cloud of non-interacting cold Fermi gas (Oppenheimer & Volkoff, 1939). However, from an observational point of view, subsequent precise measurements of NS mass from binary systems have revealed maximum masses that exceed the theoretical predictions (Hulse & Taylor, 1975; Anderson et al., 1990; Wolszczan, 1991; Burgay et al., 2003), and it suggests that the interactions between particles in the core cannot be ignored. Due to the poorly known and uncertain composition and interactions in the core of NSs, as well as the EOS of NS, the maximum mass $M_{\text{max}}$ of a NS remains highly uncertain, ranging from $1.46\,M_{\odot}$ to $2.48\,M_{\odot}$ for different EOSs (Haensel et al., 2007). So far, the maximum mass of NS that we have accurately observed is ${2.08}_{-0.07}^{+0.07}\ {M}_{\odot}$ (68.3% credibility) from PSR J0740+6620 determined by relativistic Shapiro delay (Fonseca et al., 2021).
@@
Several previous studies have attempted to constrain the EOS of neutron stars, such as through the observed internal plateau together with the collapse time in the X-ray emission of short-duration gamma-ray bursts (GRBs) (Lasky et al., 2014; Lü et al., 2015), gravitational-wave (GW) radiation from hypermassive NSs (HMNS)(Shibata, 2005), GW radiation from supramassive NSs (Lan et al., 2020), GW radiation from the merger of NS binaries (Takami et al., 2014), and the threshold mass for prompt black hole formation from double NS mergers (Bauswein et al., 2020). However, the constraints from these results are not yet sufficient. Thus, understanding the EOS of NSs remains an open question in astrophysics and particle physics (Lattimer, 2021).
@@
On the other hand, by considering the NS as a good resonator, numerous quasi-periodic oscillations (QPOs) modes of the NS would be excited. Based on their restoring forces, the oscillation modes can be categorized into radial and different non-radial modes (such as fundamental-mode, pressure-mode, gravity-mode, Rossby-mode, and shear-mode) (McDermott et al., 1988; Stergioulas, 2003). Watts & Strohmayer (2007) discovered quasi-periodic oscillations (QPOs) with frequencies ranging from tens of hertz up to a few kilohertz by analyzing the X-ray light curves of SGR 0526-66, SGR 1900+14, and SGR 1806-20. The discovery of QPOs in the three cases is believed to be the first direct detection of magnetar oscillations, and it is naturally interpreted as crustal shear modes triggered by starquakes which are associated with giant flares (GFs) (Duncan, 1998; Piro, 2005; Watts & Strohmayer, 2007; Lee, 2007; Samuelsson & Andersson, 2007; Sotani et al., 2007, 2012; Steiner & Watts, 2009). The oscillations are strongly dependent on the density distribution of the stellar interior and the shear modulus within the elastic region. By comparing the observed frequencies with theoretical models, precise constraints can be placed on the properties of ultra-dense matter which cannot be replicated to check in terrestrial laboratories (Lattimer & Prakash, 2001; Sotani et al., 2012).
@@
Recently, Chirenti et al. (2023) claimed that kilohertz quasi-periodic signals were identified in two short-duration GRBs (931101B and 910711) from BATSE data (see Table 1). They proposed that the physical origin of these two GRBs may originate from the merger of binary NSs, and such a merger results in a hot remnant, such as an HMNS. If the QPOs of those two short-duration GRBs are indeed from the quasi-periodic modulation of gamma-rays caused by the global oscillations, then it may provide a new opportunity to explore the properties of NSs by employing seismology. In this paper, we consider two potential mechanisms which are possible to produce the observed high-frequency QPOs in two short-duration GRBs. One is torsional oscillation with crustal shear modes in cold NSs (Section 2), and the other one is radial oscillation in the hot merger remnants by considering the effect of temperature (Section 3). Then, we constrain the EOS of neutron stars by adopting the observed high-frequency QPOs in those two short-duration GRBs (Section 4). Conclusions are drawn in Section 5, with some additional discussions.
@@ 2 Torsional oscillation
Soft gamma-ray repeaters (SGRs) are young, slowly spinning magnetars, and strong flaring activities exhibiting durations from milliseconds to seconds with QPOs can be produced during the active outburst phase (Duncan & Thompson, 1992). Because of the intense $\gamma-$ray luminosity and the spectral characteristics, the emission of GFs and short-duration GRBs are quite similar, and GFs of magnetars were also detected as disguised short-duration GRBs from nearby galaxies. Recently, two short GRBs (200415A and 051103A) have been interpreted as giant magnetar flares (Yang et al., 2020; Svinkin et al., 2021; Roberts et al., 2021). If this is the case, the GFs with QPOs are driven by elastic shear modes in the crust of a cold NS which is believed to consist of a fluid core with a radius $r_{\text{c}}$, and an elastic crust with a thickness of $\sim$ 1 km accounts for a mass of  $\sim\,0.01\ {M}_{\odot}$. As strong magnetic fields and/or rotational effects evolve with time, the stresses will build up in the crust. When these stresses reach a critical threshold, the crust will eventually fracture to trigger oscillations. Due to the velocity field of torsional oscillations (i.e., the restoring force supported by shear force) is divergence-free with no radial component, it is not related to the varying density and pressure of the star (Schumaker & Thorne, 1983). Thus, it is more easily excited with a fundamental frequency of $\sim$ 30 Hz after a crustquake of a magnetar with a strong magnetic field (Schumaker & Thorne, 1983; McDermott et al., 1988; Duncan, 1998).
One needs to note that Chirenti et al. (2023) excluded an SGR origin, primarily based on energetics estimates from the redshift $z$ inferred using distances to galaxies with moderately active star formation, which are consistent with the burst localizations. However, this scenario cannot be entirely ruled out as a possibility in a low-redshift context. In this section, we hypothesize that the observed short-duration GRBs with high-frequency QPOs (GRB 931103B and GRB 910711) are caused by GFs of magnetars at low redshift. Specifically, we consider $z<0.015$, corresponding to a maximum isotropic energy $E_{\text{iso}}=5.3\times 10^{46}$ erg in the GF sample (see below for details). Within this assumption, we discuss the torsional oscillations of NSs and adopt the high-frequency QPOs observed in these bursts to constrain the EOS of NSs.
@@ 2.1 Physical model
Thorne & Campolattaro (1967) first derived the equations governing the non-radial oscillation of neutron stars in General Relativity. Similarly, we consider a spherical neutron star under the hypothesis of a non-rotating and non-magnetic field. The marked feature of the torsional oscillation mode is that it does not involve radial displacement nor induce any change in internal density and pressure. Consequently, within the frame of spherical coordinates ($r$, $\theta$ and $\phi$), the defined crystalline matter element moves only along the angle $\phi$. The magnitude of proper displacement in the $\phi$-direction of a material element during oscillation can be denoted as $\mathrm{d}l$, which can be expressed as
@@
Here, $b_{\ell}(\theta)$ is related to the Legendre polynomial $P_{\ell}(\cos\theta)$, and it is the only $\theta$-dependent function that governs the angular behavior of displacement. Hence, $\ell$ is the angular index that determines the angular nodes of the oscillation. The comprehensive mathematical framework for torsional oscillations of non-rotating stars in the context of General Relativity was established by Schumaker & Thorne (1983). The function $Y(r)$ governs the radial behavior of the oscillation and can be formulated as (Schumaker & Thorne, 1983; Sotani et al., 2012)
$Y^{\prime\prime}+\left(\frac{4}{r}+\Phi^{\prime}-\Lambda^{\prime}+\frac{\mu^{\prime}}{\mu}\right)Y^{\prime}+\left[\frac{\rho+P/c^{2}}{\mu}\omega^{2}\mathrm{e}^{-2\Phi}-\frac{(\ell+2)(\ell-1)}{r^{2}}\right]\mathrm{e}^{2\Lambda}Y=0.$
@@
Finally, by giving the mass and EOS of a neutron star, the spherically symmetric density $\rho$ and pressure $P$ of the neutron star can be solved by the Tolman-Oppenheimer-Volkoff (TOV) equation. Then, the eigen-frequency $\omega$ can be obtained by solving Eq. (2.1), supplemented by the boundary conditions $Y^{\prime}(r_{\text{c}})=0$ and $Y^{\prime}(R)=0$ at the core-crust interface and the surface of NS, respectively.
@@
In general, the oscillation of a neutron star has an infinite number of discrete eigen-frequency solutions. The different modes of torsional oscillation can be labeled as ${}_{\ell}t_{n}$, which correspond to the frequency ${}_{\ell}\omega_{n}=2\pi_{\ell}\nu_{n}$. Here, the indices $n=0,1,2,...$ and $\ell=2,3,4,\ldots$ represent the nodes of oscillation in radial and angular directions, respectively. The specific excitation of oscillation modes is related to the properties of the crust and the processes of coupling and damping. However, some oscillation modes may be excited but not detected. The reason may be related to the modulation modes of the gamma-ray light curve, which remain unclear.
2.2 Constraint on the EOS of NS
@@
For the fundamental oscillations without radial nodes ($n=0$), the momentum transfer of shear perturbations in the $\theta$-direction is significantly greater than that of in the $r$-direction. If this is the case, it will result in the formation of standing waves during the propagation (with a typical length-scale of $\sim\pi R$) of shear perturbations along the $\theta$-direction. In contrast, the length-scale of propagation in the $r$-direction is  $\sim\,0.1R$ (Gabler et al., 2012). Consequently, this process takes a longer time to result in lower oscillation frequencies ${}_{\ell}\nu_{0}\sim 20-100$ Hz which are much lower than that of observed QPOs frequencies in the two short-duration GRBs 931103B and 910711 (Chirenti et al., 2023). Hence, it is believed that the origin of QPO frequencies observed in those two short-duration GRBs is not likely to relate to fundamental oscillations. On the other hand, by considering the eigen-frequency from ordinary oscillations with $n\geq 1$ (one or more radial nodes) as a function of mass, one can obtain the ‘fine splitting’ structure, which is similar to a spectrum. The frequencies change slightly with increasing values of $\ell$ compared to the change with $n$. Therefore, we adopt $\ell=2$ to replace the other modes with $\ell>2$. Initially, we consider only cold-catalyzed (zero-temperature), non-magnetic, non-rotating spherically symmetric models. The given mass of a NS is strongly dependent on the frequency spectrum. In this case, we adopt a 0.5% error in the frequency spectrum lines, which may be caused by the greater complexity of the actual neutron star (e.g., temperature, magnetic fields, rotation, and non-spherically symmetric features), but our knowledge about these complexities is limited.
@@
Then, we investigate the relationship between torsional oscillation frequency and mass for different given EOSs of NSs (see Fig. 1 and Fig. 2). In our calculations, we selected six unified EOSs which are suitable for the condition of zero-temperature: GM1 (Glendenning & Moszkowski, 1991), TM1 (Sugahara & Toki, 1994), NL3 (Lalazissis et al., 1997), APR (Akmal et al., 1998), SLy4 (Douchin & Haensel, 2001), and DDME2 (Lalazissis et al., 2005). In Fig. 1 and Fig. 2, the horizontal dashed lines and their shadows correspond to the frequencies and error ranges of the two QPOs listed in Table. 1, respectively. The vertical dashed lines and their shadows indicate the ranges of mass which are capable of producing torsional oscillations at those frequencies. Two high-frequency QPOs in the same GRB correspond to two ranges of mass, and it is expected that the ranges of mass for those two high-frequency can overlap within the margin of error.
@@
By considering the effects of the magnetic field, rotation, and non-spherical symmetry of NS, but limited knowledge about these factors, we adopt a 0.5% error in the frequency spectrum lines. For example, the magnetic field confined to the crust may affect the frequencies of various torsional modes. One can make a rough estimate of such an effect of the magnetic field $B$ by using Newtonian physics, which is expressed as (Duncan, 1998; Messios et al., 2001)
@@
where ${}_{\ell}\nu_{n}^{(0)}$ is the frequency in the non-magnetic case. Based on the numerical simulation of torsional and magneto-elastic oscillations of a realistic magnetar model, Gabler et al. (2011) found that a $\sim 10^{14}$ G magnetic field can achieve a similar damping time in the order of ten milliseconds. For given the value of magnetic field of $\sim 4\times 10^{14}$ G, it will result in a correction closer to 0.5%. This allows us to believe that the 0.5% error in the frequency spectrum we adopt is reasonable. On the other hand, the non-measured redshift ($z$) of those two short-duration GRBs also introduces significant uncertainty. The observed frequency of QPO is directly related to the frequency in the rest-frame, namely:
@@
The frequency in the rest-frame is higher than that of frequency in the observed frame with a factor ($1+z$). However, the observed maximum flux of GRB 910711 and GRB 931101B are $1.5\times 10^{-4}$ $\text{erg cm}^{-2}~{}\text{s}^{-1}$ and $2.6\times 10^{-5}$ $\text{erg cm}^{-2}~{}\text{s}^{-1}$, respectively. On the contrary, the maximum isotropic energy $E_{\text{iso}}$ in the GFs sample can only reach to $5.3\times 10^{46}$ erg (Burns et al., 2021). By assuming that the isotropic energy can reach to the maximum isotropic energy of GFs from SGR, one can derive $z<0.015$ which is corresponding to a frequency error less than $0.15\%$. However, it should be emphasized that the error of 0.5% is only a simple assumption for an uncertainty originating from various complexities. In more extensive error range cases, the ranges of mass required to generate a lower-frequency QPO may overlap simultaneously with multiple ranges of mass (i.e. the intersection of the higher-frequency QPO and spectrum lines with different $n$). This results in two overlapping areas in the mass range, or even the range of masses required for lower frequency QPO may be wholly covered. Thus, in this case, the mass range is roughly determined by the QPO with the lower frequency.
For given six EOSs of NS in Fig. 1 and Fig. 2, it is found that all selected EOSs exhibit significant overlapping mass ranges for both GRB 931101B and GRB 910711. It suggests that a NS with a maximum mass of  $\sim\,1.38\ M_{\odot}$ (GM1),  $\sim\,1.88\ M_{\odot}$ (TM1),  $\sim\,2.01\ M_{\odot}$ (NL3),  $\sim\,1.70\ M_{\odot}$ (APR),  $\sim\,1.73\ M_{\odot}$ (SLy4), and  $\sim\,2.28\ M_{\odot}$ (DDME2) would satisfy the criteria to be considered as a potential source for the QPOs observed in GRB 931101B, and a maximum mass of  $\sim\,1.79\ M_{\odot}$ (GM1),  $\sim\,2.15\ M_{\odot}$ (TM1),  $\sim\,2.52\ M_{\odot}$ (NL3),  $\sim\,2.09\ M_{\odot}$ (APR),  $\sim\,2.01\ M_{\odot}$ (SLy4), and  $\sim\,2.48\ M_{\odot}$ (DDME2) for GRB 910711 (see Table. 2). In other words, we cannot constrain the EOS of NS via the high-frequency QPOs in these two short-duration GBRs if we believe that they originated from the GFs of magnetar.

### v2

# Constraining the Nuclear Equation of State of a Neutron Star via High-frequency Quasi-periodic Oscillation in Short Gamma-Ray Bursts

@@ Abstract
The determination of the equation of state (EOS) of a neutron star (NS) and its maximum mass is very important for understanding the formation and properties of NSs under extreme conditions, but they remain open questions. Short-duration gamma-ray bursts (GRBs) are believed to originate from the merger of binary NSs or giant flares (GFs) of soft gamma repeaters (SGRs). Recently, the high-frequency quasi-periodic oscillations (QPOs) have been claimed to be identified from two short GRBs (GRB 931101B and GRB 910711). In this paper, we propose that the observed high-frequency QPOs in these two short GRBs result from torsional oscillations in the GFs of SGRs associated with cold NSs, or from radial oscillations of hypermassive NSs as the hot remnants of binary NS mergers, and then to constrain the EOS of NSs. For torsional oscillations, the six selected EOSs (TM1, NL3, APR, SLy4, DDME2, and GM1) of NSs suitable for the zero-temperature condition exhibit significant overlap in mass ranges, suggesting that we cannot constrain the EOS of NSs. For radial oscillations, the six selected EOSs (IUF, TM1, TMA, FSG, BHBLp, and NL3) of NSs suitable for the high-temperature condition cannot be ruled out when redshift is considered. However, it is found that the EOS can only be constrained if the redshift and temperature of the remnant can be measured.
@@ 1 Introduction
Neutron stars (NSs) are one of the most intriguing compact objects in astrophysical studies, due to their extreme properties. The maximum mass and equation of state (EOS) of NSs play an important role in understanding their structure and properties (Shapiro & Teukolsky, 1983). The first estimate of the maximum mass of an NS ($M_{\text{max}}$) for a static equilibrium solution in early studies was $M_{\text{max}}\sim\,0.72\ M_{\odot}$ Me by simplifying the NS as a cloud of noninteracting cold Fermi gas (Oppenheimer & Volkoff, 1939). However, from an observational point of view, subsequent precise measurements of NS mass from binary systems have revealed maximum masses that exceed the theoretical predictions (Hulse & Taylor, 1975; Anderson et al., 1990; Wolszczan, 1991; Burgay et al., 2003), and this suggests that the interactions between particles in the core cannot be ignored. Due to the poorly known and uncertain composition and interactions in the core of NSs, as well as the EOS of NSs, the maximum mass $M_{\text{max}}$ of an NS remains highly uncertain, ranging from $1.46\,M_{\odot}$ to $2.48\,M_{\odot}$ for different EOSs (Haensel et al., 2007). So far, the maximum mass of an NS that we have accurately observed is ${2.08}_{-0.07}^{+0.07}\ {M}_{\odot}$ (68.3% credibility) from PSR J0740+6620, determined by relativistic Shapiro delay (Fonseca et al., 2021).
@@
Several previous studies have attempted to constrain the EOS of NSs, such as through the observed internal plateau together with the collapse time in the X-ray emission of short-duration gamma-ray bursts (GRBs; Lasky et al. (2014); Lü et al. (2015)) , gravitational-wave (GW) radiation from hypermassive NSs (HMNS; Shibata (2005)), GW radiation from supramassive NSs (Lan et al., 2020), GW radiation from the merger of NS binaries (Takami et al., 2014), and the threshold mass for prompt black hole (BH) formation from double NS mergers (Bauswein et al., 2020). However, the constraints from these results are not yet sufficient. Thus, understanding the EOS of NSs remains an open question in astrophysics and particle physics (Lattimer, 2021).
@@
On the other hand, by considering the NS as a good resonator, numerous quasi-periodic oscillation (QPOs) modes of the NS would be excited. Based on their restoring forces, the oscillation modes can be categorized into radial and different nonradial modes (such as fundamental mode, pressure mode, gravity mode, Rossby mode, and shear mode; McDermott et al. (1988); Stergioulas (2003)). Watts & Strohmayer (2007) discovered QPOs with frequencies ranging from tens of Hz up to a few kHz by analyzing the X-ray light curves of SGR 0526–66, SGR 1900+14, and SGR 1806–20. The discovery of QPOs in the three cases is believed to be the first direct detection of magnetar oscillations, and it is naturally interpreted as crustal shear modes triggered by starquakes, which are associated with giant flares (GFs; Duncan (1998); Piro (2005); Watts & Strohmayer (2007); Lee (2007); Samuelsson & Andersson (2007); Sotani et al. (2007, 2012); Steiner & Watts (2009)). The oscillations are strongly dependent on the density distribution of the stellar interior and the shear modulus within the elastic region. By comparing the observed frequencies with theoretical models, precise constraints can be placed on the properties of ultradense matter that cannot be replicated to check in terrestrial laboratories (Lattimer & Prakash, 2001; Sotani et al., 2012).
@@
Recently, Chirenti et al. (2023) claimed that kHz quasi-periodic signals were identified in two short-duration GRBs (GRB 931101B and GRB 910711) from BATSE data (see Table 1). They proposed that the physical origin of these two GRBs may originate from the merger of binary NSs, and such a merger results in a hot remnant, such as an HMNS. If the QPOs of those two short-duration GRBs are indeed from the quasi-periodic modulation of gamma rays caused by the global oscillations, then it may provide a new opportunity to explore the properties of NSs by employing seismology. In this paper, we consider two potential mechanisms that may produce the observed high-frequency QPOs in two short-duration GRBs. One is torsional oscillation with crustal shear modes in cold NSs (Section 2), and the other one is radial oscillation in the hot merger remnants by considering the effect of temperature (Section 3). Then, we constrain the EOS of neutron stars (NSs) by adopting the observed high-frequency QPOs in those two short-duration GRBs (Section 3.2). Conclusions are drawn in Section 4, with some additional discussions.
@@ 2 Torsional Oscillation
Soft gamma repeaters (SGRs) are young, slowly spinning magnetars, and strong flaring activities exhibiting durations from milliseconds to seconds with QPOs can be produced during the active outburst phase (Duncan & Thompson, 1992). Because of the intense $\gamma$-ray luminosity and the spectral characteristics, the emission of GFs and that of short-duration GRBs are quite similar, and GFs of magnetars were also detected as disguised short-duration GRBs from nearby galaxies. Recently, two short GRBs (GRB 200415A and GRB 051103A) have been interpreted as giant magnetar flares (Yang et al., 2020; Svinkin et al., 2021; Roberts et al., 2021). If this is the case, the GFs with QPOs are driven by elastic shear modes in the crust of a cold NS that is believed to consist of a fluid core with a radius $r_{\text{c}}$, and an elastic crust with a thickness of $\sim$ 1 km accounts for a mass of  $\sim\,0.01\ {M}_{\odot}$. As strong magnetic fields and/or rotational effects evolve with time, the stresses will build up in the crust. When these stresses reach a critical threshold, the crust will eventually fracture to trigger oscillations. Because the velocity field of torsional oscillations (i.e., the restoring force supported by shear force) is divergence-free with no radial component, it is not related to the varying density and pressure of the star (Schumaker & Thorne, 1983). Thus, it is more easily excited with a fundamental frequency of $\sim$ 30 Hz after a crustquake of a magnetar with a strong magnetic field (Schumaker & Thorne, 1983; McDermott et al., 1988; Duncan, 1998).
One needs to note that Chirenti et al. (2023) excluded an SGR origin, primarily based on energetics estimates from the redshift $z$ inferred using distances to galaxies with moderately active star formation, which are consistent with the burst localizations. However, this scenario cannot be entirely ruled out as a possibility in a low-redshift context. In this section, we hypothesize that the observed short-duration GRBs with high frequency QPOs (GRB 931103B and GRB 910711) are caused by GFs of magnetars at low redshift. Specifically, we consider $z<0.015$, corresponding to a maximum isotropic energy $E_{\text{iso}}=5.3\times 10^{46}$ erg in the GF sample (see below for details). Within this assumption, we discuss the torsional oscillations of NSs and adopt the high-frequency QPOs observed in these bursts to constrain the EOS of NSs.
@@ 2.1 Physical Model
Thorne & Campolattaro (1967) first derived the equations governing the nonradial oscillation of NSs in general relativity. Similarly, we consider a spherical NS under the hypothesis of a nonrotating and nonmagnetic field. The marked feature of the torsional oscillation mode is that it does not involve radial displacement or induce any change in internal density and pressure. Consequently, within the frame of spherical coordinates ($r$, $\theta$ and $\phi$), the defined crystalline matter element moves only along the angle $\phi$. The magnitude of proper displacement in the $\phi$-direction of a material element during oscillation can be denoted as $\mathrm{d}l$, which can be expressed as
@@
Here $b_{\ell}(\theta)$ is related to the Legendre polynomial $P_{\ell}(\cos\theta)$, and it is the only $\theta$-dependent function that governs the angular behavior of displacement. Hence, $\ell$ is the angular index that determines the angular nodes of the oscillation. The comprehensive mathematical framework for torsional oscillations of nonrotating stars in the context of general relativity was established by Schumaker & Thorne (1983). The function $Y(r)$ governs the radial behavior of the oscillation and can be formulated as (Schumaker & Thorne, 1983; Sotani et al., 2012)
$Y^{\prime\prime}+\left(\frac{4}{r}+\Phi^{\prime}-\Lambda^{\prime}+\frac{\mu^{\prime}}{\mu}\right)Y^{\prime}+\left[\frac{\rho+P/c^{2}}{\mu}\omega^{2}\mathrm{e}^{-2\Phi}-\frac{(\ell+2)(\ell-1)}{r^{2}}\right]\mathrm{e}^{2\Lambda}Y=0.$ (2)
@@
Finally, by giving the mass and EOS of an NS, the spherically symmetric density $\rho$ and pressure $P$ of the NS can be solved by the Tolman–Oppenheimer–Volkoff equation. Then, the eigenfrequency $\omega$ can be obtained by solving Equation (2), supplemented by the boundary conditions $Y^{\prime}(r_{\text{c}})=0$ and $Y^{\prime}(R)=0$ at the core–crust interface and the surface of the NS, respectively.
@@
In general, the oscillation of an NS has an infinite number of discrete eigenfrequency solutions. The different modes of torsional oscillation can be labeled as ${}_{\ell}t_{n}$, which correspond to the frequency ${}_{\ell}\omega_{n}=2\pi_{\ell}\nu_{n}$. Here the indices $n=0,1,2,...$ and $\ell=2,3,4,\ldots$ represent the nodes of oscillation in radial and angular directions, respectively. The specific excitation of oscillation modes is related to the properties of the crust and the processes of coupling and damping. However, some oscillation modes may be excited but not detected. The reason may be related to the modulation modes of the gamma-ray light curve, which remain unclear.
2.2 Constraint on the EOS of NSs
@@
For the fundamental oscillations without radial nodes ($n=0$), the momentum transfer of shear perturbations in the $\theta$-direction is significantly greater than that in the $r$-direction. If this is the case, it will result in the formation of standing waves during the propagation (with a typical length scale of $\sim\pi R$) of shear perturbations along the $\theta$-direction. In contrast, the length scale of propagation in the $r$-direction is  $\sim\,0.1R$ (Gabler et al., 2012). Consequently, this process takes a longer time to result in lower oscillation frequencies ${}_{\ell}\nu_{0}\sim 20-100$ Hz, which are much lower than those of observed QPOs in the two shortduration GRBs GRB 931103B and GRB 910711 (Chirenti et al., 2023). Hence, it is believed that the origin of QPO frequencies observed in those two short-duration GRBs is not likely to relate to fundamental oscillations. On the other hand, by considering the eigenfrequency from ordinary oscillations with $n\geq 1$ (one or more radial nodes) as a function of mass, one can obtain the “fine splitting” structure, which is similar to a spectrum. The frequencies change slightly with increasing values of $\ell$ compared to the change with $n$. Therefore, we adopt $\ell=2$ to replace the other modes with $\ell>2$. Initially, we consider only cold-catalyzed (zero-temperature), nonmagnetic, nonrotating spherically symmetric models. The given mass of an NS is strongly dependent on the frequency spectrum. In this case, we adopt a 0.5% error in the frequency spectrum lines, which may be caused by the greater complexity of the actual NS (e.g., temperature, magnetic fields, rotation, and nonspherically symmetric features), but our knowledge about these complexities is limited.
@@
Then, we investigate the relationship between torsional oscillation frequency and mass for different given EOSs of NSs (see Figures 1 and 2). In our calculations, we selected six unified EOSs that are suitable for the condition of zero temperature: GM1 (Glendenning & Moszkowski, 1991), TM1 (Sugahara & Toki, 1994), NL3 (Lalazissis et al., 1997), APR (Akmal et al., 1998), SLy4 (Douchin & Haensel, 2001), and DDME2 (Lalazissis et al., 2005). In Figures 1 and 2, the horizontal dashed lines and shading correspond to the frequencies and error ranges of the two QPOs listed in Table 1, respectively. The vertical dashed lines and shading indicate the ranges of mass that are capable of producing torsional oscillations at those frequencies. Two high-frequency QPOs in the same GRB correspond to two ranges of mass, and it is expected that the ranges of mass for those two high-frequency QPOs can overlap within the margin of error.
@@
By considering the effects of the magnetic field, rotation, and nonspherical symmetry of NS, but with limited knowledge about these factors, we adopt a 0.5% error in the frequency spectrum lines. For example, the magnetic field confined to the crust may affect the frequencies of various torsional modes. One can make a rough estimate of such an effect of the magnetic field B by using Newtonian physics, which is expressed as (Duncan, 1998; Messios et al., 2001)
@@
where ${}_{\ell}\nu_{n}^{(0)}$ is the frequency in the nonmagnetic case. Based on the numerical simulation of torsional and magneto-elastic oscillations of a realistic magnetar model, Gabler et al. (2011) found that a $\sim 10^{14}$ G magnetic field can achieve a similar damping time on the order of tens of milliseconds. For the given value of magnetic field of $\sim 4\times 10^{14}$ G, it will result in a correction closer to 0.5%. This allows us to believe that the 0.5% error in the frequency spectrum we adopt is reasonable. On the other hand, the nonmeasured redshift ($z$) of those two short-duration GRBs also introduces significant uncertainty. The observed frequency of QPOs is directly related to the frequency in the rest frame, namely,
@@
The frequency in the rest frame is higher than that in the observed frame with a factor ($1+z$). However, the observed maximum fluxes of GRB 910711 and GRB 931101B are $1.5\times 10^{-4}$ $\text{erg cm}^{-2}~{}\text{s}^{-1}$ and $2.6\times 10^{-5}$ $\text{erg cm}^{-2}~{}\text{s}^{-1}$, respectively. On the contrary, the maximum isotropic energy Eiso in the GF sample can only reach to $5.3\times 10^{46}$ erg (Burns et al., 2021). By assuming that the isotropic energy can reach to the maximum isotropic energy of GFs from SGRs, one can derive $z<0.015$, which corresponds to a frequency error of less than 0.15%. However, it should be emphasized that the error of 0.5% is only a simple assumption for an uncertainty originating from various complexities. In more extensive error range cases, the ranges of mass required to generate a lower-frequency QPO may overlap simultaneously with multiple ranges of mass (i.e., the intersection of the higher-frequency QPO and spectrum lines with different $n$). This results in two overlapping areas in the mass range, or even the range of masses required for lower-frequency QPOs may be wholly covered. Thus, in this case, the mass range is roughly determined by the QPO with the lower frequency.
For the six given EOSs of NSs in Figures 1 and 2, it is found that all selected EOSs exhibit significant overlapping mass ranges for both GRB 931101B and GRB 910711. This suggests that an NS with a maximum mass of  $\sim\,1.38\ M_{\odot}$ (GM1),  $\sim\,1.88\ M_{\odot}$ (TM1),  $\sim\,2.01\ M_{\odot}$ (NL3),  $\sim\,1.70\ M_{\odot}$ (APR),  $\sim\,1.73\ M_{\odot}$ (SLy4), and  $\sim\,2.28\ M_{\odot}$ (DDME2) would satisfy the criteria to be considered as a potential source for the QPOs observed in GRB 931101B, and similarly for a maximum mass of  $\sim\,1.79\ M_{\odot}$ (GM1),  $\sim\,2.15\ M_{\odot}$ (TM1),  $\sim\,2.52\ M_{\odot}$ (NL3),  $\sim\,2.09\ M_{\odot}$ (APR),  $\sim\,2.01\ M_{\odot}$ (SLy4), and  $\sim\,2.48\ M_{\odot}$ (DDME2) for GRB 910711 (see Table 2). In other words, we cannot constrain the EOS of NSs via the high-frequency QPOs in these two short-duration GBRs if we believe that they originated from the GFs of magnetars.